OpenStudy (anonymous):

if pa=6 and pe=4, find be

OpenStudy (raden):

do you have a diagram ?

OpenStudy (anonymous):

https://media.glynlyon.com/g_geo_2013/7/group172.gif

OpenStudy (raden):

let be is x. PE * PB = PA^2 4 * (4+x) = 6^2 solve for x

OpenStudy (anonymous):

ok so how would i do it i realy stink at this i dropped out in 2010 and just started back to night class this week and dont ever remember it being this hard

OpenStudy (raden):

4 * (4+x) = 6^2 4 * (4+x) = 36 both sides divided by 4, get (4+x) = 9 subtracted by 4 on both sides, get x = 5